# Random Turán Problems for Hypergraph Expansions

### v1

@@ 5 Applying our Results
Here we apply the Lifting Theorems we have developed to give short proofs of Theorems 1.2 and 1.4 which give tight bounds for the random Turán problem for expansions of $K_{2,t}$ and theta graphs $\theta_{a,b}$. The details of the related result 1.3 for $K_{s,t}^{(r)}$ with general $s$ is deferred to an arXiv only appendix since the computations are cumbersome and since we prove stronger bounds for $s\geq 3$ in forthcoming work [16]. Our proof relies on 1.7, for which we need the following.
@@
By 3.3 it suffices to construct a tight $3$-tree whose 2-shadow contains a copy of $\theta_{a,b}$. Let $u$ and $v$ be the two end points of $\theta_{a,b}$, let $P_{i}$, $1\leq i\leq a$ be the $a$ paths, and let $w_{i,j}$, $11\leq j\leq b-$ 1, be the internal vertices of $P_{i}$, and for notational convenience we let $w_{i,0}=u$ and $w_{i,b}=v$ for all $i$. To build our tight tree, we begin by adding all of the edges of the form $\{w_{i,0},w_{i,1},w_{i,b}\}$ for all $i$, and iteratively given then we have added all edges of the form $\{w_{i,j-1},w_{i,j},w_{i,b}\}$ for some $21\leq j\leq b-$ 2 we add all edges of the form $\{w_{i,j},w_{i,j+1},w_{i,b}\}$. It is not difficult to see that this is a tight tree (since each edge we add adds the new vertex $w_{i,j+1}$) and that its 2-shadow contains the edge $w_{i,j},w_{i,j+1}$ for all $1\leq i\leq a$ and $11\leq j\leq b-$ 1 (with the $j=b-1$ case coming from the final edges $\{w_{i,b-2},w_{i,b-1},w_{i,b}\}$, and hence a copy of $\theta_{a,b}$. ∎
We now show that $\theta^{(r)}_{a,b}$ has balanced balanced supersaturation whenever $\theta_{a,b}$ does with $a\geq 3$.
@@
We begin by using 1.7 to show that $\theta_{a,b}^{(3)}$ is balanced, and for this we need to check the hypothesis of 1.7. By Lemma 5.1, $\theta^{(3)}_{a,b}$ is a spanning subgraph of a tight 3-tree. Property (a) holds by hypothesis, and it is easy to check that $\gamma_{2}$ and $\tau_{2}$ satisfy properties (b) and (c). Thus Theorem 1.7 guarantees the existence of some sufficiently large constant $C_{3}$ such that the following holds. Let $M_{3}(n)=C_{3}n^{2}$, let $\gamma_{3}(n)=C_{2}C_{3}\log n$, and let
@@
Letting $\tau_{3}^{\prime}(n,m)=nn^{1+\frac{a-1}{ab-1}}(\log n)^{C_{3}}$, the above shows that $\theta^{(3)}_{a,b}$ is $(M_{3},\gamma_{3},\tau_{3}^{\prime})$-balanced and a spanning subgraph of a tight $3$-tree. The desired random Turán bound then follows for all $r\geq 3$ by 1.8. ∎
To utilize 5.2, we make use of the following known balanced supersaturation results (rephrased in the language of our paper).
@@
Theorems 1.2 and 1.4 now follow immediately from 5.2 together with these balanced supersaturation results (and the fact that $K_{2,t}=\theta_{t,2}$).

### v2

@@ 6 Proofs of Main Results
We now apply our Lifting Theorems to prove our results. We begin by proving our main result 2.1 for large expansions in Section 6.1, after which we give short proofs of Theorems 2.2 and 2.4 giving tight bounds for the random Turán problem for expansions of $K_{2,t}$ and theta graphs $\theta_{a,b}$ for smaller values of $r$ in Section 6.2. The details of the related result 2.3 for $K_{s,t}^{(r)}$ with general $s$ is deferred to an arXiv-only appendix as the computations are cumbersome and we plan to prove even stronger bounds for $s\geq 3$ in forthcoming work [28].
6.1 Proof of 2.1
The statement of 2.1 involves Sidorenko hypergraphs, for which we make a related definition.
Definition 5.
We say that a hypergraph is $(C,\alpha)$-Sidorenko if every $n$-vertex $r$-graph $H$ with $|H|\geq Cn^{\alpha}$ contains at least $C^{-1}(|H|n^{-r})^{|F|}n^{v(F)}$ copies of $F$.
It is well known that being $(C,\alpha)$-Sidorenko for some $\alpha<r$ is equivalent to being Sidorenko, though for our purposes, we will only need the following easy-to-prove direction of this relationship.
Lemma 6.1.
If $F$ is a non-empty $r$-graph which is Sidorenko, then there exists some $C>0$ such that $F$ is $(C,r-\frac{1}{|F|})$-Sidorenko.
Proof.
Let $C$ be a sufficiently large constant and let $H$ be an $r$-graph with $|H|\geq Cn^{r-1/|F|}$. Because $F$ is Sidorenko, the number of homomorphisms from $F$ to $H$ is at least $(|H|n^{-r})^{|F|}n^{v(F)}\geq C^{|F|}n^{v(F)-1}$. Since the number of non-injective homomorphisms from $F$ to $H$ is trivially at most $O(n^{v(F)-1})$, we conclude that, given $C$ is sufficiently large, at least a half of the homomorphisms from $F$ to $H$ are injective, i.e. there are at least $\frac{1}{2}(|H|n^{-r})^{|F|}n^{v(F)}$ genuine copies of $F$ in $H$. Thus taking $C$ sufficiently large gives the result. ∎
We next state a hypergraph analog of a balanced supersaturation result for general graphs due to Jiang and Longbrake [13, Theorem 3.7]. For this, given an $r$-graph $F$ with $|F|\geq 3$, we define its proper $r$-density $d^{*}_{r}(F)$ by
$d^{*}_{r}(F):=\max_{F^{\prime}\subsetneq F,\ |F^{\prime}|\geq 2}\frac{|F^{\prime}|-1}{v(F^{\prime})-r}.$
Theorem 6.2.
Let $F$ be a non-empty $r$-partite $r$-graph with $|F|\geq 3$. If there exist positive reals $C$ and $\alpha>r-\frac{1}{d_{r}(F)}$ such that $F$ is $(C,\alpha)$-Sidorenko, then there exists a constant $C^{\prime}=C^{\prime}(F)>0$ such that $F$ is $(M,\gamma,\tau)$-balanced where $M(n)=C^{\prime}n^{r-\frac{1}{|F|}}$, $\gamma(n)=C^{\prime}$, and
$\tau(n,m)=C^{\prime}\max\left\{n^{r-\frac{v(F)-r}{|F|-1}},~{}m^{1-\frac{1}{d_{r}^{*}(F)(r-\alpha)}}n^{\frac{\alpha}{d^{*}_{r}(F)(r-\alpha)}}\right\}.$
The proof of 6.2 is an entirely straightforward generalization of [13, Theorem 3.7] from graphs to hypergraphs, and as such we defer the exact details of this argument to an arXiv-only appendix.
The following lemma provides some inequalities that will be useful in the upcoming proofs. For this, we recall from the introduction that we say an $r$-graph $F$ with $|F|\geq 3$ is strictly $r$-balanced if $d_{r}(F)=\frac{|F|-1}{v(F)-r}$ and if (in terms of the new notation above) we have $d_{r}(F)>d^{*}_{r}(F)$.
Lemma 6.3.
If $F$ is a strictly balanced $r_{0}$-graph with $\Delta(F)\geq 2$ and $r\geq|F|^{2}v(F)r_{0}$, then
  • (i)
    $r_{0}-\frac{v(F)-r_{0}}{|F|-1}>0,$
  • (ii)
    $\frac{r_{0}|F|-1}{d^{*}_{r_{0}}(F)}+\left(1-\frac{|F|}{d_{r_{0}}^{*}(F)}\right)\left(\frac{r-r_{0}}{r-1}\right)>0,$
  • (iii)
    $\left(r_{0}-\frac{1}{|F|}\right)\frac{r-1}{r_{0}-1}-\frac{r-r_{0}}{r_{0}-1}<r-1,$
  • (iv)
    $r_{0}-\frac{v(F)-r_{0}}{|F|-1}\geq\frac{r_{0}|F|-1}{d_{r_{0}}^{*}(F)}+\left(r_{0}-\frac{r_{0}-1}{r-1}\right)\left(1-\frac{|F|}{d_{r_{0}}^{*}(F)}\right).$
Proof.
Part (i) is equivalent to saying $v(F)<r_{0}|F|$. And indeed, since $F$ is balanced and has at least one edge, $F$ has no isolated vertices, implying that $v(F)\leq r_{0}|F|$ with equality only if $F$ is a matching. However, $F$ can not be a matching since $\Delta(F)\geq 2$, so this inequality must be strict.
For part (ii), since $10<\frac{r-r_{0}}{r-1}<$ 1, the lefthand side of the inequality is larger than
$\frac{r_{0}|F|-1}{d^{*}_{r_{0}}(F)}-\frac{|F|}{d^{*}_{r_{0}}(F)}\geq\frac{|F|-1}{d^{*}_{r_{0}}(F)}>0.$
For part (iii), moving all monomials with $r$ to the righthand side gives
$\frac{1}{|F|(r_{0}-1)}+1<\frac{r}{|F|(r_{0}-1)},$
which is equivalent to $r>|F|(r_{0}-1)+1$. This clearly holds since $r\geq|F|^{2}v(F)r_{0}$.
For part (iv), by moving all of the terms in the inequality depending on $r$ to one side and using that $\frac{v(F)-r_{0}}{|F|-1}=\frac{1}{d_{r_{0}}(F)}$ since $F$ is balanced, we find that the stated inequality is equivalent to
$\frac{1}{d_{r_{0}}^{*}(F)}-\frac{1}{d_{r_{0}}(F)}\geq\frac{r_{0}-1}{r-1}\left(\frac{|F|}{d^{*}_{r_{0}}(F)}-1\right),$
which after multiplying by $d_{r_{0}}(F)d_{r_{0}}(F)$ on both sides gives
$d_{r_{0}}(F)-d_{r_{0}}^{*}(F)\geq\frac{r_{0}-1}{r-1}(|F|-d^{*}_{r_{0}}(F))d_{r_{0}}(F).$
We claim that
$d_{r_{0}}(F)-d^{*}_{r_{0}}(F)\geq\frac{1}{\left(v(F)-r_{0}\right)^{2}}.$
Indeed, this follows from the elementary fact that if $a/b$ and $c/d$ are rationals with $a/b>c/d$ and $1\leq b,d\leq t$, then
$\frac{a}{b}-\frac{c}{d}=\frac{ad-bc}{bd}\geq\frac{1}{t^{2}}.$
On the other hand,
$\frac{r_{0}-1}{r-1}\cdot\left(|F|-d^{*}_{r_{0}}(F)\right)\cdot d_{r_{0}}(F)\leq\frac{r_{0}}{r}\cdot|F|\cdot\frac{|F|}{v(F)-r_{0}}.$
Thus it suffices to check that
$\frac{1}{\left(v(F)-r_{0}\right)^{2}}\geq\frac{r_{0}}{r}\cdot|F|\cdot\frac{|F|}{v(F)-r_{0}},$
which is true since $r\geq|F|^{2}v(F)r_{0}$.
Using the Lifting Theorems (2.7 and 2.8) together with 6.1 and 6.2, we are now ready to show that $F^{(r)}$ has optimal balanced supersaturation when $r$ is sufficiently large.
Theorem 6.4.
Let $F$ be a strictly $r_{0}$-balanced $r_{0}$-graph with $\Delta(F)\geq 2$. If $F$ is Sidorenko, then for all $r>|F|^{2}v(F)r_{0}$, there exists a constant $C>0$ such that $F^{(r)}$ is $(M,\gamma,\tau)$-balanced where $M(n)=Cn^{r-1}$, $\gamma=(\log n)^{C}$, and
$\tau(n,m)=n^{r-\frac{1}{d_{r}(F^{(r)})}}(\log n)^{C}.$
Proof.
Since $F$ is Sidorenko, we know that $F$ is $r_{0}$-partite, and by 6.1, we know that there exists a constant $C_{1}$ such that $F$ is $(C_{1},r_{0}-\frac{1}{|F|})$-Sidorenko. Note that $d_{r_{0}}(F)=\frac{|F|-1}{v(F)-r_{0}}<|F|$, which implies $r_{0}-\frac{1}{|F|}>r_{0}-\frac{1}{d_{r_{0}}(F)}$. Thus by 6.2, there exists a constant $C_{r_{0}}>0$ such that $F$ is $(M_{r_{0}},\gamma_{r_{0}},\tau_{r_{0}})$-balanced where $M_{r_{0}}(n)=C_{r_{0}}n^{r_{0}-\frac{1}{|F|}}$, $\gamma_{r_{0}}(n)=C_{r_{0}}$, and
$\tau_{r_{0}}(n,m)=C_{r_{0}}\max\left\{n^{r_{0}-\frac{v(F)-r_{0}}{|F|-1}},~{}m^{1-\frac{|F|}{d_{r_{0}}^{*}(F)}}n^{\frac{r_{0}|F|-1}{d^{*}_{r_{0}}(F)}}\right\}.$
Let $r_{1}=|F|^{2}v(F)r_{0}$. We want to apply 2.7 to “lift” the balanced supersaturation results for $F$ to $F^{(r_{1})}$. To this end, we need to check that conditions (a), (b), (c) and (d) in 2.7 hold. Conditions (a), (b) and (c) (that $F$ is $(M_{r_{0}},\gamma_{r_{0}},\tau_{r_{0}})$-balanced, that $M_{r_{0}}(n)n^{-\frac{r_{1}-r_{0}}{r_{1}-1}}$ is non-decreasing, and that $\gamma_{r_{0}}(n)$ is non-decreasing) are straightforward to verify. To check condition (d) (that $\tau_{r_{0}}(n,m)$ is non-increasing with $m$ and that $\tau_{r_{0}}(nx,mx^{\frac{r_{1}-r_{0}}{r_{1}-1}})$ is non-decreasing with $x$), note that the exponent of $m$ in $\tau_{r_{0}}(n,m)$ is either 0 or $01-\frac{|F|}{d^{*}_{r_{0}}(F)}<$ 0, and that the exponent of $x$ in $\tau_{r_{0}}(nx,mx^{\frac{r_{1}-r_{0}}{r_{1}-1}})$ is either
[…]
which are both positive by 6.3. Thus condition (d) holds. By 2.7 we know that there exists a sufficiently large constant $C_{r_{1}}$ such that $F^{(r_{1})}$ is $(M_{r_{1}},\gamma_{r_{1}},\tau_{r_{1}})$-balanced where
$M_{r_{1}}(n)=\max\left\{M_{r_{0}}(n)^{\frac{r_{1}-1}{r_{0}-1}}n^{-\frac{r_{1}-r_{0}}{r_{0}-1}},~{}n^{r_{1}-1}\right\}(\log n)^{C_{r_{1}}}=n^{r_{1}-1}(\log n)^{C_{r_{1}}},$
where here the second equality used 6.3(iii),
$\gamma_{r_{1}}(n)=\gamma_{r_{0}}(n)\left(\log n\right)^{C^{\prime}}\leq\left(\log n\right)^{C_{r_{1}}},$
and
$\tau_{r_{1}}(n,m)$ $=\tau_{r_{0}}\left(n,n^{\frac{r_{1}-r_{0}}{r_{1}-1}}m^{\frac{r_{0}-1}{r_{1}-1}}(\log n)^{-C^{\prime}}\right)(\log n)^{C^{\prime}}$
$\leq\max\left\{n^{r_{0}-\frac{v(F)-r_{0}}{|F|-1}},~{}\left(n^{\frac{r_{1}-r_{0}}{r_{1}-1}}m^{\frac{r_{0}-1}{r_{1}-1}}\right)^{\left(1-\frac{|F|}{d_{r_{0}}^{*}(F)}\right)}n^{\frac{r_{0}|F|-1}{d^{*}_{r_{0}}(F)}}\right\}(\log n)^{C_{r_{1}}}$
$\leq\max\left\{n^{r_{0}-\frac{v(F)-r_{0}}{|F|-1}},~{}n^{\left(\frac{r_{1}-r_{0}}{r_{1}-1}+r_{0}-1\right)\left(1-\frac{|F|}{d_{r_{0}}^{*}(F)}\right)}n^{\frac{r_{0}|F|-1}{d^{*}_{r_{0}}(F)}}\right\}(\log n)^{C_{r_{1}}}$
$=n^{r_{0}-\frac{v(F)-r_{0}}{|F|-1}}(\log n)^{C_{r_{1}}},$
where here the second inequality uses the fact that $m\geq M_{r_{1}}(n)\geq n^{r_{1}-1}$ and that $01-\frac{|F|}{d^{*}_{r_{0}}(F)}<$ 0, and the last equality is implied by 6.3(iv).
For $r>r_{1}$, we want to get rid of the logarithmic factor in $M(n)$ by using 2.8 on $F^{(r-1)}$. Note that $r>r_{1}>v(F)$, and thus $F^{(r)}$ is a spanning subgraph of a tight $r$-tree by 4.3 (since the $r_{0}$-shadow of a single $v(F)$-edge contains $F$ as a spanning subgraph). Suppose that we have inductively proven that there exists a constant $C_{r-1}$ such that $F^{(r-1)}$ is $(M_{r-1},\gamma_{r-1},\tau_{r-1})$-balanced where $M_{r-1}(n)=n^{r-2}(\log n)^{C_{r-1}}$, $\gamma_{r-1}(n)=(\log n)^{C_{r-1}}$, and
$\tau_{r-1}(n,m)=n^{r_{0}-\frac{v(F)-r_{0}}{|F|-1}}(\log n)^{C_{r-1}},$
noting that we have proved this for the case $r-1=r_{1}$. It is straightforward to check that conditions (a), (b) and (c) in 2.8 hold. Let $C_{r}$ be a sufficiently large constant and let $M_{r}(n)=C_{r}n^{r-1}$, $\gamma_{r}(n)=\gamma_{r-1}(n)\cdot C_{r}(\log n)$, and let
$A=A(n,m)=\frac{m}{M_{r-1}(n)\cdot C_{r}\log n}=\frac{m}{n^{r-2}\cdot C_{r}(\log n)^{C_{r-1}+1}}.$
Clearly, condition (d) in 2.8 holds. Thus by 2.8, $F^{(r)}$ is $(M_{r},\gamma_{r},\tau_{r})$-balanced where there exists a sufficiently large constant $C$ such that
$\tau_{r}(n,m)=\max\left\{A(n,m)^{-\frac{1}{d_{r}(F^{(r)})}}\cdot m,~{}\tau_{r-1}\left(n,~{}\frac{m}{A(n,m)\cdot C_{r}\log n}\right)\cdot C_{r}\log n\right\}\leq n^{r-\frac{1}{d_{r}(F^{(r)})}}(\log n)^{C},$
where here the last inequality comes from the definition of $\tau_{r-1}$ together with the fact from 4.2 that
$r_{0}-\frac{v(F)-r_{0}}{|F|-1}=r_{0}-\frac{1}{d_{r_{0}}(F)}=r-\frac{1}{d_{r}(F^{(r)})},$
and that (using $m\geq n^{r-1}$)
$A(n,m)^{-\frac{1}{d_{r}(F^{(r)})}}\cdot m=m^{1-\frac{1}{d_{r}(F^{(r)})}}n^{\frac{r-2}{d_{r}(F^{(r)})}+o(1)}\leq n^{r-1-\frac{1}{d_{r}(F^{(r)})}+o(1)}.$
This completes the proof. ∎
As an aside, we note that it is perhaps a little surprising that although the general result of Jiang and Longbrake 6.2 typically does not give optimal balanced supersaturation results for most $r_{0}$-graphs $F$, the proof of 6.4 shows that our Lifting Theorems are powerful enough to translate these potentially suboptimal bounds into optimal bounds for high enough expansions.
Finally, we will need the following result from [29, Theorem 1.4] in order to prove the “moreover” portion of 2.1.
Theorem 6.5 ([29]).
If $F$ is an $r$-graph with $|F|\geq 2$ and $\frac{v(F)-r}{|F|-1}<r$ which is not Sidorenko, then there exists some $\epsilon>0$ such that for any $p=p(n)\geq n^{-\frac{v(F)-r}{|F|-1}}$, we have a.a.s.
$\mathrm{ex}(G_{n,p}^{r},F)\geq n^{r-\frac{v(F)-r}{|F|-1}-o(1)}(pn^{\frac{v(F)-r}{|F|-1}})^{\epsilon}.$
The exact statement of [29, Theorem 1.4] in fact replaces the $\epsilon$ here with a more precise quantity $\frac{s(F)}{e(F)-1+s(F)}$ where $s(F)$ measures how “far” from Sidorenko $F$ is, but all that we need here is $F$ being non-Sidorenko implies $s(F)>0$. With this, we now have all the pieces needed to prove 2.1.
Proof of 2.1.
The first part of the theorem regarding Sidorenko $F$ follows immediately from 2.9 and 6.4. For the moreover portion, let $F$ be as in the hypothesis of the theorem such that $F^{(r)}$ is not Sidorenko for all $r\geq r_{0}$. By the definition of the expansion $F^{(r)}$ and $F$ being balanced, we have for all $r\geq r_{0}$ that
$r-\frac{v(F^{(r))}-r}{|F^{(r)}|-1}=r-\frac{v(F)+(r-r_{0})|F|-r}{|F|-1}=r_{0}-\frac{v(F)-r_{0}}{|F|-1}=r_{0}-\frac{1}{d_{r_{0}}(F)},$
and by 6.3(i) this quantity is positive. This together with 6.5 implies that for all $r\geq r_{0}$, there exists some $\epsilon>0$ such that for all $\delta>0$ and $p=n^{-r+r_{0}-\frac{1}{d_{r_{0}}(F)}+\delta}=n^{-r+\frac{v(F^{(r)})-r}{|F^{(r)}|-1}+\delta}$, we have a.a.s.
$\mathrm{ex}(G_{n,p}^{r},F^{(r)})\geq n^{r_{0}-\frac{1}{d_{r_{0}}(F)}+\epsilon\delta-o(1)}.$
This bound is strictly stronger than what is claimed in 2.1 for Sidorenko hypergraphs whenever $10<\delta<$ 1, proving that such bounds can not extend to any $F^{(r)}$. ∎
Before moving on, we emphasize that the bound of $r>|F|^{2}v(F)r_{0}$ obtained in 2.1 can be improved upon significantly provided more information about the hypergraph $F$ is known. Notably, one can obtain better bounds here provided one has better bounds on either (a) the difference $d_{r_{0}}(F)-d_{r_{0}}^{*}(F)$, (b) the smallest $r_{1}$ such that $F^{(r_{1}+1)}$ is contained in a spanning subgraph of a tight $(r_{1}+1)$-tree, or (c) the smallest $\alpha$ such that $F$ is $(C,\alpha)$-Sidorenko for some $C$. In particular, our current proof uses the most pessimistic estimates possible for each of these quantities, so there is quite a bit of room for substantial improvement for specific choices of $F$.
More precisely, if one uses more precise bounds on $d_{r_{0}}(F)-d^{*}_{r_{0}}(F)$, $r_{1}$ and $\alpha$, then a more involved calculation shows that the tight bounds of 2.1 would hold provided
$r\geq r_{0}-\frac{1}{d_{r_{0}}(F)}+\frac{(r_{0}-1)\left((r_{1}-r_{0})d_{r_{0}}(F)+1\right)(\frac{1}{r_{0}-\alpha}-d^{*}_{r_{0}}(F))}{(r_{1}-1)(d_{r_{0}}(F)-d^{*}_{r_{0}}(F))}.$
In particular, when $F=C_{2\ell}$, we can use $r_{1}=r_{0}=2$, $d_{2}(C_{2\ell})=\frac{2\ell-1}{2\ell-2}$, $d^{*}_{2}(C_{2\ell})=1$, and $\alpha=1+\frac{1}{\ell}$ to obtain tight bounds for $r\geq 4$, recovering the known bounds for even cycles as stated in 1.1. However, we emphasize that one can often obtain better bounds for $r$ than even this optimized version of our result whenever more precise balanced supersaturation results are known beyond the general results of Jiang and Longbrake. In particular, for the case $F=K_{2,t}$, this optimized result above gives tight bounds only for $r\geq t+2$, while (as we show below) the true answer is $r\geq 3$ for $t\geq 3$.
6.2 Proofs of 2.2 and 2.4
Our proof for theta graphs relies on 2.8, for which we need the following.
@@
By 4.3 it suffices to construct a tight $3$-tree whose 2-shadow contains a copy of $\theta_{a,b}$. Let $u$ and $v$ be the two end points of $\theta_{a,b}$, let $P_{i}$, $1\leq i\leq a$ be the $a$ paths, and let $w_{i,j}$, $11\leq j\leq b-$ 1, be the internal vertices of $P_{i}$, and for notational convenience we let $w_{i,0}=u$ and $w_{i,b}=v$ for all $i$. To build our tight tree, we begin by adding all of the edges of the form $\{w_{i,0},w_{i,1},w_{i,b}\}$ for all $i$, and iteratively given then we have added all edges of the form $\{w_{i,j-1},w_{i,j},w_{i,b}\}$ for some $21\leq j\leq b-$ 2 we add all edges of the form $\{w_{i,j},w_{i,j+1},w_{i,b}\}$. It is not difficult to see that this is a tight tree (since each edge we add adds the new vertex $w_{i,j+1}$) and that its 2-shadow contains the edge $w_{i,j},w_{i,j+1}$ for all $1\leq i\leq a$ and $11\leq j\leq b-$ 1 (with the $j=b-1$ case coming from the final edges $\{w_{i,b-2},w_{i,b-1},w_{i,b}\}$, and hence a copy of $\theta_{a,b}$. ∎
We now show that $\theta^{(r)}_{a,b}$ has optimal balanced supersaturation whenever $\theta_{a,b}$ does with $a\geq 3$.
@@
We begin by using 2.8 to show that $\theta_{a,b}^{(3)}$ is balanced, and for this we need to check the hypothesis of 2.8. By Lemma 6.6, $\theta^{(3)}_{a,b}$ is a spanning subgraph of a tight 3-tree. Property (a) holds by hypothesis, and it is easy to check that $\gamma_{2}$ and $\tau_{2}$ satisfy properties (b) and (c). Thus Theorem 2.8 guarantees the existence of some sufficiently large constant $C_{3}$ such that the following holds. Let $M_{3}(n)=C_{3}n^{2}$, let $\gamma_{3}(n)=C_{2}C_{3}\log n$, and let
@@
Letting $\tau_{3}^{\prime}(n,m)=nn^{1+\frac{a-1}{ab-1}}(\log n)^{C_{3}}$, the above shows that $\theta^{(3)}_{a,b}$ is $(M_{3},\gamma_{3},\tau_{3}^{\prime})$-balanced and a spanning subgraph of a tight $3$-tree. The desired random Turán bound then follows for all $r\geq 3$ by 2.9. ∎
To utilize 6.7, we make use of the following known balanced supersaturation results (rephrased in the language of our paper).
@@
Theorems 2.2 and 2.4 now follow immediately from 6.7 together with these balanced supersaturation results (and the fact that $K_{2,t}=\theta_{t,2}$).